Nhat is the equivalent resistance Rab? 4Ω 2Ω b 8Ω 4Ω Ο 6Ω Ο 4 Ω Ο 5Ω Ο 3 Ω
Answer #1

Req= (4+2)||(8+4)

Req=6||12

Req=(6*12)/(6+12)

Req=72/18

Req=4 ohm

So, second option is correct.